1.) You have 10 identical resistors. a.) How would you make the largest equivalent resistance out of these resistors? b.) How would you make an equivalent resistance of the lowest value out of these resistors? (resistance is 220kΩ)
Answer #1

For series combination Req = R1 + R2 + R3 + .... R10 =) Req = 10R

For parallel combination 1/Req = 1/R1 + 1/R2 + 1/R3 ..... + 1/R10

=) Req = R/10

a.) For larger equivalent resistance , combination should be series . Req = 10×220 = 2200k.ohm

b.) For smaller equivalent resistance , combination should be in parallel . Req = 220/10 = 22k.ohm